# MCGMark: An Encodable and Robust Online Watermark for LLM-Generated Malicious Code

### v1

@@ 4.3.1 Code Quality Enhancement with Outliers
However, outliers can affect the accuracy of watermark detection. During detection, we can only analyze the code and cannot obtain the real-time probability distribution of the vocabulary. This limitation aligns with real-world scenarios. To address this, we include error-correction bits in the watermark to recover the watermark information. When outliers impact the watermark information, MCGMark sets the error-correction bit to $1$; otherwise, it is set to 00. Once the watermark detection bits are fully embedded, the error-correction bits are also generated. Subsequently, the error-correction bits are embedded into the code. During the embedding of error-correction bits, the watermark is not influenced by outliers. This ensures accurate watermark detection while reducing the impact of the watermark on the embedded elements by more than $50\%$.
4.3.2 Outliers Detection
MCGMark employs the Inter-Quartile Range (IQR) method for outlier detection, based on the boxplot approach [57]. The IQR method is a robust statistical technique for describing data dispersion, measuring the distribution range of the middle $50\%$ of data by calculating the difference between the third and first quartiles of a dataset [58]. This method is particularly suitable for MCGMark due to its robustness to extreme values and its effectiveness in identifying outliers in non-normally distributed data, which is common in token probability distributions [59, 60]. MCGMark defines the upper whisker of the boxplot as:
$\small F_{\text{upper}}=Q_{3}+S\cdot IQR=(S+1)\cdot Q_{3}-S\cdot Q_{1}$ (1)
where $Q_{1}$ is the lower quartile, $Q_{3}$ is the upper quartile, $IQR=Q_{3}-Q_{1}$, and $S$ is a scaling factor. This approach allows us to identify and appropriately handle tokens that are crucial for maintaining code generation quality.
@@ 4.3.2 Outliers Detection
Input: Prompt: $R$, watermark: $W$, the number of new tokens: $L$, threshold for $P_{\text{dis }}$: $Thr\_P_{\text{dis }}$, a hash key $H$.
1:  for $l=0,1,\cdots,L$ do
2:     Obtain $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$ and $P_{l}=\left\{P_{0}^{l},P_{1}^{l},\cdots,P_{n}^{l}\right\}$.
3:     Calculate the outliers of $P_{l}$ with Equ. (1).
4:     if $F_{\text{upper }}\neq\emptyset$ then
5:        Case 1: $|F_{\text{upper }}|=1$,$D^{{}^{\prime}}=D\cup F_{\text{upper }}$.
6:        Case 2: $|F_{\text{upper }}|\geq 2$, $D^{{}^{\prime}}=D\cup F_{\text{upper }}[0:\lceil\frac{F_{\text{upper }}}{2}\rceil]$ with $H$.
7:     end if
8:     Sample $v_{i}^{(l)}$ from $V$.
9:     if $v_{i}^{(l)}\in F_{\text{upper}}~{}\&~{}v_{i}^{(l)}\notin D$ then
10:        Setting the watermark’s error-correction bit to $1$.
11:     end if
12:  end for
4.4 Enhancing Watermark Robustness
In this process, we provide a detailed description of MCGMark’s robustness scheme, designed based on the structure of malicious code. This scheme enables watermarks to withstand simple code modifications attempted by malicious developers. We delineate the adversarial scenarios and present a comprehensive overview of the design process.
4.4.1 Adversarial Scenarios
Malicious developers utilize LLM to generate malicious code. When these developers become aware of the potential presence of a watermark in the code, they may modify it. To avoid compromising the code’s execution, malicious developers tend to modify easily changeable elements, especially for less experienced “script kiddies.” Consequently, if the watermark is placed on easily modifiable elements, such as variable names, the watermark becomes highly susceptible to being rendered ineffective.
4.4.2 The Overall of Watermark Robustness Enhancement
This step presents the robustness enhancement scheme designed by MCGMark to address adversarial scenarios. Enhancing the robustness of watermarking in LLM-generated code involves skipping embedding watermarks in parts of the code that are easily modified or removed. However, since code generation and watermark embedding need to be synchronized, MCGMark must decide online whether to skip watermark embedding for the next token. To achieve this, MCGMark divides the process into two steps.
  1. 1.
    MCGMark determines which elements in the code should be skipped for watermark embedding. This is done by defining watermark skipping rules based on code structure, code cloning, and observations from malicious code instances.
  2. 2.
@@ 4.4.2 The Overall of Watermark Robustness Enhancement
It is worth noting that due to significant differences in code structure among different programming languages, and since Python is currently one of the most commonly used languages by malicious developers [61], MCGMark focuses solely on Python language.
4.4.3 Watermark Skipping Rules
In this step, we describe the element selection criteria of MCGMark for watermark skipping. To design watermark skipping rules, it is crucial to identify which elements in Python code are easily modifiable without affecting code usability. Our analysis focuses on three key perspectives:
  • (Code Structure.) Python primarily consists of the following elements [62]: indentation, keywords, identifiers, statements, expressions, functions, objects, object methods, types, numbers, operators, comments, exception handling, input/output, and blank lines. Previous research indicates that modifications to certain elements have minimal impact on code execution quality: identifiers (including variable names, function names, and class names), comments, output statements, numerical values, and blank lines [63, 64].
  • (Code Clone.) Code clone focusing on identifying instances of code plagiarism using similarity metrics [65, 66]. And it’s classification typically into four levels: exact, lexical, syntactical, and semantic. Among these, exact clones are identical except for whitespace, layout, and comments. Lexical clones differ in identifiers but maintain the exact same structure. Syntactical clones involve the addition, modification, or deletion of statements while preserving the overall structure. Semantic clones, on the other hand, refer to code segments that achieve the same functionality but use different syntax. Recent studies indicate that detecting complete and lexical similarity is feasible [67, 68, 69, 70, 71, 72, 73], suggesting that modifying elements like whitespace, layout, comments, and identifiers is more manageable [74, 75, 76].
@@ 4.4.3 Watermark Skipping Rules
In summary, the watermark is not embedded in code elements listed in Table LABEL:tab:EazyModi.
4.4.4 Watermark Skipping Pattern
In this step, we design watermark skip patterns for MCGMark. The watermark embedding skip algorithm in MCGMark is based on watermark skipping rules. Since LLM-generated tokens are irreversible and the watermark embedding process is synchronized with the token generation, MCGMark cannot wait for the model to generate elements in Table LABEL:tab:EazyModi before skipping the watermark. The embedding modifies the distribution of $P_{l}=\left\{P_{0}^{l},P_{1}^{l},\cdots,P_{n}^{l}\right\}(l\in[0,L])$, influencing the LLM’s decision-making. Therefore, watermark embedding must be controlled before generating elements in Table LABEL:tab:EazyModi. Thus, MCGMark decides on watermark embedding for the next token based on the already generated code, considering that vocabulary tokens are irregular and may not correspond directly to code elements.
@@ 4.4.4 Watermark Skipping Pattern
  • (Pattern 6.) When one Pattern is in effect, another Pattern may be triggered, leading to conflicts in Pattern decisions. In such cases, if any Pattern is already active, subsequent Patterns will not be triggered. However, if conditions in $v_{i}^{(l)}$ satisfy the triggering criteria of two Patterns simultaneously, only the first Pattern in sequence will be triggered.
  • (Pattern 7.) Once all watermark bits have been added, but not all tokens have been generated, MCGMark continue iterating the embedding process to further strengthen the watermark.
4.4.5 Watermark Skipping Process
Algorithm 2 Enhancing the Robustness of LLM Online Code Watermark via Code Structure and Syntax
Input: Existing tokens: $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$ , watermark to be embedded $w_{X}$, set $\widehat{A,B,C,D}$.
1:  if $l\leq L$ then
@@ 4.4.5 Watermark Skipping Process
3:     if not $\mathcal{LOCK}$ then
4:        if $v_{i}^{(l)}\in\left\{\widehat{A}\cup\widehat{B}\cup\widehat{C}\cup\widehat{D}\right\}$ then
@@
7:           $V=V$, break.
8:        else
9:           if $X\leq x$ then
10:              Take set $D$ or $V\cap D$ corresponding to $w_{X}$, to $V$, $X=X+1$.
11:           else
@@
13:           end if
14:        end if
15:     end if
16:     Based on the effectiveness of Pattern to determine $\mathcal{LOCK}\Leftarrow 0$.
17:     $V=V$, break.
18:  end if
In this step, we describe MCGMark’s watermark skip decision process based on the established skip rules and patterns. Algorithm 2 delineates the execution process of watermark skipping. After obtaining the sequence $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$, MCGMark first verify if $l\leq L$, where $L$ represents the maximum output token limit of the LLM. If this condition is met, MCGMark proceeds with the watermarking process; otherwise, MCGMark terminates. MCGMark then evaluates $v_{i}^{(l)}$ against several conditions: if it consists solely of whitespace characters, MCGMark triggers Pattern 5; MCGMark checks for any active patterns based on Pattern 6 which would preclude watermarking the next token; and MCGMark determines if $v_{i}^{(l)}\in\left\{\widehat{A}\cup\widehat{B}\cup\widehat{C}\cup\widehat{D}\right\}$, which would trigger the corresponding Pattern (1, 2, 3, or 4) along with Pattern 6 to prevent multi-pattern conflicts. If no pattern is triggered, MCGMark selects a word from vocabulary $\mathbb{A}$ or $\mathbb{B}$ based on the watermark information. Upon complete embedding of all watermark information, MCGMark triggers Pattern 7 to iterate and incorporate additional watermark data. Algorithm 2 ensures judicious watermark embedding while preserving code integrity and functionality, accounting for various code elements and potential pattern conflicts.
In summary, after these steps, the watermarks embedded by MCGMark are imperceptible and encode the code generator’s personal information. This approach simultaneously ensures code quality and enhances robustness by leveraging code structure.
4.5 Design and Detection of Watermark
4.5.1 Watermark Design
This step delineates the watermark prompt design for MCGMark. MCGMark employs a dual-component watermark design comprising Detection Bits and Error Correction Bits. Detection Bits primarily encode user information for traceability, while Error Correction Bits facilitate the recovery of potentially erroneous information in the Detection Bits. Although Algorithm 1 ensures LLM code generation quality, setting a low outlier threshold (small $S$ value) may result in excessive outliers, potentially impacting the LLM’s word selection. For instance, the LLM might be compelled to select words from vocabulary $\mathbb{B}$ instead of $\mathbb{A}$ as dictated by the watermark information bit, due to outlier presence. This scenario could lead to errors in specific watermark bits. To mitigate this issue, MCGMark introduces Error Correction Bits to restore information in the Detection Bits and generates watermarks in the Detection Bits without outlier influence. Furthermore, Detection Bits and Error Correction Bits are designed with equal length. This design effectively addresses the trade-off between maintaining code quality and preserving watermark integrity, enhancing the overall robustness of MCGMark’s watermarking strategy.
4.5.2 Watermark Detection
This step describes the lightweight watermark detection process for MCGMark. In detecting watermarks within code, MCGMark requires the code to be inspected, the model’s vocabulary, tokenizer, Algorithm 1, and Algorithm 2. There is no need to load an additional LLM. Given the malicious code, MCGMark first tokenizes the code using the tokenizer. Next, it applies the seven patterns and Algorithm 2 to remove elements where watermark embedding can be skipped, resulting in a sequence of code elements. Then, the hash value $H$ in Algorithm 1 is used to partition the vocabulary. The code element sequence is then traversed, and elements are categorized into the corresponding vocabulary parts, producing a sequence of 00s and $1$s. MCGMark subsequently segments this sequence according to the watermark length. Each segment provides detection bits and error-correction bits, which are used to obtain the user’s identity information using the following formula:
$\text{Detection Bits}\oplus(\text{1}~{}\&~{}\text{Error Correction Bits})~{}.$ (2)
Since our watermark is embedded in a round-robin manner, multiple segments of the watermark can sometimes be extracted from the code. When multiple rounds of watermarking are possible, consistent results from at least two rounds can help preliminarily identify the malicious code generator, thus enhancing fault tolerance.
5 Experiments
We evaluate MCGMark in this section based on the watermark design requirements analyzed in Section 2.3 with MCGTest. Specifically, we focus on the following research questions:

### v2

@@ 4.3. Ensuring Watermarked Code Quality
However, outliers can affect the accuracy of watermark detection. During detection, we can only analyze the code and cannot obtain the real-time probability distribution of the vocabulary. This limitation aligns with real-world scenarios. To address this, we include error-correction bits in the watermark to recover the watermark information. When outliers impact the watermark information, MCGMark sets the error-correction bit to $1$; otherwise, it is set to 00. Once the watermark detection bits are fully embedded, the error-correction bits are also generated. Subsequently, the error-correction bits are embedded into the code. During the embedding of error-correction bits, the watermark is not influenced by outliers.
Outliers Detection. MCGMark uses the Inter-Quartile Range (IQR) method, based on boxplot statistics (Bondarenko et al., 2024), to detect outliers in token probability distributions. IQR measures the spread of the middle $50\%$ of data by computing the difference between the third ($Q_{3}$) and first ($Q_{1}$) quartiles (Vinutha et al., 2018). This method is well-suited for MCGMark due to its robustness to extreme values and its effectiveness in non-normally distributed data (Powell et al., 2023; Yang et al., 2019). We define the upper whisker as Equ. (1), where $S$ is a scaling factor.
(1) $F_{\text{upper}}=Q_{3}+S\cdot IQR=(S+1)\cdot Q_{3}-S\cdot Q_{1}.$
By detecting outliers in the LLM’s vocabulary during code generation, we are able to identify and handle tokens that are critical for maintaining code quality. This mechanism helps avoid selecting low-quality tokens that could compromise the readability or functionality of the code during watermark embedding.
@@
4.4. Enhancing Watermark Robustness
In this process, we provide a detailed description of MCGMark’s robustness scheme, designed based on the code structure. This scheme enables watermarks to withstand typical code modifications attempted by malicious developers. We delineate the adversarial scenarios and present a comprehensive overview of the design process.
Adversarial Scenarios. Once malicious developers become aware of the possible presence of a watermark, they may attempt to tamper with the code. To avoid breaking its functionality, such developers, especially those with limited experience, tend to modify easily changeable elements, such as comments or variable names. Consequently, if the watermark is placed on easily modifiable elements, such as variable names, the watermark becomes highly susceptible to being rendered ineffective.
The Overall of Watermark Robustness Enhancement. This step introduces the robustness enhancement strategy of MCGMark to address adversarial scenarios. To improve the resilience of watermarking in LLM-generated code, MCGMark avoids embedding watermarks in code regions that are easily modified or removed. However, since watermark embedding must be synchronized with code generation, MCGMark needs to make real-time decisions on whether to embed a watermark for each token. To enable this, the embedding process is divided into two stages.
  1. (1)
    MCGMark identifies code elements to be excluded from watermark embedding by defining skipping rules grounded in code structure, code cloning patterns, and insights from real-world malicious code.
  2. (2)
@@ 4.4. Enhancing Watermark Robustness
It is worth noting that due to significant differences in code structure among different programming languages, and since Python is currently one of the most commonly used languages by malicious developers (medium, 2024), MCGMark focuses solely on Python language.
Watermark Skipping Rules. In this step, we describe the element selection criteria of MCGMark for watermark skipping. To design watermark skipping rules, it is crucial to identify which elements in Python code are easily modifiable without affecting code usability. Our analysis focuses on three key perspectives:
  • (Code Structure.) Python primarily consists of the following elements (Schwarz et al., 2020): indentation, keywords, identifiers, statements, expressions, functions, objects, object methods, types, numbers, operators, comments, exception handling, input/output, and blank lines. Previous research indicates that modifications to certain elements have minimal impact on code execution quality: identifiers (including variable names, function names, and class names), comments, output statements, numerical values, and blank lines (Zhang et al., 2023b; Funabiki et al., 2022).
  • (Code Clone.) Code clone detection focuses on identifying plagiarized code using similarity metrics (Xu et al., 2024; Li et al., 2024a), typically categorized into four levels: exact, lexical, syntactical, and semantic. Exact clones differ only in whitespace, layout, or comments; lexical clones use different identifiers but retain structure; syntactical clones modify statements while preserving structure; semantic clones perform the same function with different syntax. Recent studies show that detecting exact and lexical clones is feasible (Dou et al., 2023; Sheneamer and Kalita, 2016; Ain et al., 2019; Min and Li Ping, 2019; Zhong et al., 2022; Lei et al., 2022; Kaur and Rattan, 2023), indicating that modifications like whitespace, layout, comments, and identifiers are relatively easy (Singh et al., 2021; Haq and Caballero, 2021; Khazaal and Asma’a, 2022).
@@
Table 1. Code Elements Excluded from Watermarking
Perspective Elements in code that are susceptible to modification
Code Structure identifiers, comments, output, numbers, blank lines
Code Clone exact clone features, lexical clone features
Malicious Code Instances comments, output, identifiers, assignments, comparisons
In summary, the watermark is not embedded in code elements listed in Table 1.
Watermark Skipping Pattern. In this step, we define watermark skipping patterns in MCGMark, guided by skipping rules. Since LLM-generated tokens are irreversible and the watermark embedding process is synchronized with the token generation, MCGMark cannot wait for the model to generate elements in Table 1 before skipping the watermark. The embedding modifies the distribution of $P_{l}=\left\{P_{0}^{l},P_{1}^{l},\cdots,P_{n}^{l}\right\}(l\in[0,L])$, influencing the LLM’s decision-making. Therefore, watermark embedding must be controlled before generating elements in Table 1. Thus, MCGMark decides on watermark embedding for the next token based on the already generated code, considering that vocabulary tokens are irregular and may not correspond directly to code elements.
Algorithm 2 Enhancing the Robustness of Watermark via Code Structure and Syntax
1:  Input: Existing tokens: $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$ , watermark to be embedded $w_{X}$, set $\widehat{A,B,C,D}$.
2:  if $l\leq L$ then
[…]
4:     if not $\mathcal{LOCK}$ then
5:        if $v_{i}^{(l)}\in\left\{\widehat{A}\cup\widehat{B}\cup\widehat{C}\cup\widehat{D}\right\}$ then
[…]
8:           $V=V$, break.
9:        else
10:           if $X\leq x$ then
11:              Take set $D$ or $V\cap D$ corresponding to $w_{X}$, to $V$, $X=X+1$.
12:           else
[…]
14:           end if
15:        end if
16:     end if
17:     Based on the effectiveness of Pattern to determine $\mathcal{LOCK}\Leftarrow 0$.
18:     $V=V$, break.
19:  end if
@@
  • (Pattern 6.) If one Pattern is active, conflicting Patterns are blocked from triggering. However, if conditions in $v_{i}^{(l)}$ satisfy the triggering criteria of two Patterns simultaneously, only the first Pattern in sequence will be triggered.
  • (Pattern 7.) Once all watermark bits are embedded, but tokens are still being generated, MCGMark continues embedding to reinforce the watermark.
It is important to note that the aforementioned rollbacks refer only to the rollback of watermark information, not the rollback of tokens generated by the LLM. When watermark information is embedded into code elements that are prone to modification or removal, the watermark needs to roll back to ensure it remains intact and avoids being altered.
Watermark Skipping Process. In this step, we describe MCGMark’s watermark skip decision process based on the established skip rules and patterns. Algorithm 2 delineates the execution process of watermark skipping. After obtaining the sequence $R^{(l)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(l)}\right\}$, MCGMark first verify if $l\leq L$, where $L$ represents the maximum output token limit of the LLM. If this condition is met, MCGMark proceeds with the watermarking process; otherwise, MCGMark terminates. MCGMark then evaluates $v_{i}^{(l)}$ against several conditions: if it consists solely of whitespace characters, MCGMark triggers Pattern 5; MCGMark checks for any active patterns based on Pattern 6 which would preclude watermarking the next token; and MCGMark determines if $v_{i}^{(l)}\in\left\{\widehat{A}\cup\widehat{B}\cup\widehat{C}\cup\widehat{D}\right\}$, which would trigger the corresponding Pattern (1, 2, 3, or 4) along with Pattern 6 to prevent multi-pattern conflicts. If no pattern is triggered, MCGMark selects a word from vocabulary $\mathbb{A}$ or $\mathbb{B}$ based on the watermark information. Upon complete embedding of all watermark information, MCGMark triggers Pattern 7 to iterate and incorporate additional watermark data. Algorithm 2 ensures judicious watermark embedding while preserving code integrity and functionality, accounting for various code elements and potential pattern conflicts.
@@
Algorithm 3 The overview of MCGMark embedding.
1:  Input: Prompt: $R$, watermark: $W$, number of tokens: $L$, threshold $Thr\_P_{\text{dis}}$, hash key $H$, set $\widehat{A,B,C,D}$.
2:  for $l=0,1,\cdots,L$ do
@@
4:        Start watermark embedding from the next token.
5:        Randomly partition the vocabulary into two parts using $H$.
6:        Call Algorithm 2 to check for skipping or rolling back watermark embedding.
7:        if embedding is required then
8:           Call Algorithm 1 to obtain outliers of the current vocabulary.
9:           If sampled outlier $\notin$ partition: set tolerance bit $=1$, else: set tolerance bit $=0$.
10:           Sample outlier and generate the token.
11:        else if rollback is required then
12:           Update $H$ to hash of the last valid token.
13:        else
14:           Do not partition the vocabulary.
15:        end if
16:     end if
17:  end for
4.5. Design and Detection of Watermark
Watermark Design. This step delineates the watermark format design for MCGMark. MCGMark employs a dual-component watermark design comprising Detection Bits and Error Correction Bits. Detection Bits primarily encode user information for traceability, while Error Correction Bits facilitate the recovery of potentially erroneous information in the Detection Bits. Although Algorithm 1 ensures LLM code generation quality, setting a low outlier threshold (small $S$ value) may result in excessive outliers, potentially impacting the LLM’s word selection. For instance, the LLM might be compelled to select words from vocabulary $\mathbb{B}$ instead of $\mathbb{A}$ as dictated by the watermark information bit, due to outlier presence. This scenario could lead to errors in specific watermark bits. To mitigate this issue, MCGMark introduces Error Correction Bits to restore information in the Detection Bits and generates watermarks in the Detection Bits without outlier influence. Furthermore, Detection Bits and Error Correction Bits are designed with equal length. This design effectively addresses the trade-off between maintaining code quality and preserving watermark integrity, enhancing the overall robustness of MCGMark’s watermarking strategy.
Algorithm 4 The overview of MCGMark detecting.
1:  Input: Code, a hash key $H$, set $\widehat{A,B,C,D}$.
2:  Tokenize the code into a sequence of tokens.
3:  Call Algorithm 2 to remove specific tokens.
4:  for each token in the sequence do
@@ 4.5. Design and Detection of Watermark
6:        Start watermark detection from the next token.
7:        for each subsequent token do
8:           Partition the vocabulary using $H$, verify token existence, and extract binary information.
9:           Update $H$ based on the last valid token.
10:        end for
11:        Split the watermark information by bit length.
12:        if bit length $<$ watermark length then
13:           Report detection failure.
14:        else
15:           If multiple rounds exist and conflict, report failure; else output watermark.
16:        end if
17:     end if
18:  end for
It is important to note that the Error Correction Bits in MCGMark differ from the traditional concept of error correction codes in network protocols. In MCGMark, the Error Correction Bits are an essential component of the watermark itself, rather than an auxiliary feature. They serve as a core functionality of the watermarking mechanism.
Watermark Detection. This step describes the lightweight watermark detection process for MCGMark. In this process, MCGMark requires the code to be detect, the LLM’s vocabulary, tokenizer, Algorithm 1, and Algorithm 2. There is no need to load any LLM. Given the malicious code, MCGMark first tokenizes the code using the tokenizer. Next, it applies the seven patterns and Algorithm 2 to remove elements where watermark embedding can be skipped, resulting in a sequence of code elements. Then, the hash value $H$ in Algorithm 1 is used to partition the vocabulary. The code element sequence is then traversed, and elements are categorized into the corresponding vocabulary parts, producing a sequence of 00s and $1$s. MCGMark subsequently segments this sequence according to the watermark length. Each segment provides detection bits and error-correction bits, which are used to obtain the user’s identity information using the following formula:
(2) $\text{Detection Bits}\oplus(\text{1}~{}\&~{}\text{Error Correction Bits})~{}.$
Due to round-robin embedding, multiple watermark segments may be extracted. Consistent results from at least two rounds help identify the malicious code generator, improving fault tolerance.
@@
5. Experiments
We evaluate MCGMark in this section based on the watermark design requirements analyzed in Section 2.3. Specifically, we focus on the following research questions: